ADNAN MUHAMMED

Grade 12,

Convert the following temperatures to the Kelvin scale. 373°C.

Convert the following temperatures to the Kelvin scale.
373°C.
 
 

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

373°C into K
373 + 273 = 646 K

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
To convert °C into K , we need to add 273.
So, 373°C into K 
= 373 + 273 = 646 K
 
Hope it helps!
Thank you ✌️